OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (2, 5).

OpenStudy (a_clan):

It's a simple problem. Your equation is y=mx+c where m=3 So, y = 3x + c Now, since (2,5) lies on this line , it must satisfy this equation. Putting x=2,y=5 in above eqn. 5=3*2 +c c= -1 Again putting back in the same eqn. y = 3x - 1 (Answer)

OpenStudy (anonymous):

thank you!